Find A quadratic Model for each set of values:
(-1,1), (1,1), (3,9)

Answer by Alan3354(69443)

Find A quadratic Model for each set of values:
(-1,1), (1,1), (3,9)
--------------
Use (x-a)*(x-b) = 0, where a & b are the 2 numbers.
For the 1st one:
(x+1)*(x-1) = 0
x^2 - 1 = 0
-------------